Use the midpoint formula to find the midpoint of a line segment with endpoints: (2, 7) and (-4, 1)
Answer by checkley77(12844)

Y=(7-1)/2=6/2=3 7-3=4 & 1+3 =4 FOR THE Y MIDPOINT
X=(2+4)/2=6/2=3 -4+3=-1 & 2-3=-1 FOR THE X MIDPOINT
(-1,4) ANS.